OpenStudy (anonymous):

Find the value of y for a given value of x, if y varies directly with x. If y = 2.52 when x = 8.4, what is y when x = 2.7?

OpenStudy (anonymous):

(2.52/8.4)*2.7=0.81

OpenStudy (opcode):

Direct variation: \(k = \frac{Y}{X}\) If y varies directly with x: \(y=kx\) \(2.52 = k(8.4) ;0.3 = k\) \(y = 0.3 * 2.7 = 0.81\)

OpenStudy (anonymous):

One of the users ask this ! Go there and get answer !:) http://openstudy.com/users/e.ali#/updates/52369c48e4b0af32a07997ba I think :)

OpenStudy (anonymous):

Any problem ?:)